OpenStudy (anonymous):

Find the sum of the three expressions and choose the correct answer. -8a^3b + 3a^2b^2 - 5 2 + 5a^2b^2 3a^3b + 7 - 9a^2b^2

OpenStudy (kropot72):

Group like terms. For example: -8a^3b + 3a^3b = ? Can you do this?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

no

OpenStudy (anonymous):

answers -5a^3b - a^2b^2 + 2 -5a^3b + 8a^2b^2 + 4 -5a^3b - a^2b^2 + 4

OpenStudy (kropot72):

Can you work out this example? -5x + 7x = ?

OpenStudy (anonymous):

-2x

OpenStudy (anonymous):

i think

OpenStudy (kropot72):

-5x + 7x = 2x Do you understand?

OpenStudy (anonymous):

y positive not negative

OpenStudy (kropot72):

Can you try this again -8a^3b + 3a^3b = ? Don't change 'a^3b', just focus on the leading numbers of each term

OpenStudy (anonymous):

ok so you get -5

OpenStudy (kropot72):

Yes, and including a^3b the answer is -5a^3b. So now we have the first term of the answer. To find the next term of the answer find this: 3a^2b^2 - 9a^2b^2 = ?

OpenStudy (anonymous):

-6

OpenStudy (kropot72):

Nearly right. But the full answer is -6a^2b^2. Now to get the second term of the answer we need to add 5a^2b^2 -6a^2b^2 + 5a^2b^2 = ?

OpenStudy (anonymous):

-1a^2b^2

OpenStudy (kropot72):

Correct! Now we just need to do the numbers to get the last term of the answer: -5 + 2 + 7 = ?

OpenStudy (anonymous):

4

OpenStudy (kropot72):

Correct! Now we just gather our three terms for the answer, remembering that -1a^2b^2 = -a^2b^2 -5a^3b -a^2b^2 +4 is the answer.

OpenStudy (anonymous):

ok thank you

OpenStudy (kropot72):

You're welcome. Good on you for perservering :)